OpenStudy (anonymous):

What is the lateral area of a cylinder whose diameter is 4 cm and whose height is three times the radius. Use 3.14 for .

OpenStudy (lgbasallote):

\[\text{LSA} = 3.14 \times r^2 \times h\]

OpenStudy (lgbasallote):

note: r = d/2

OpenStudy (lgbasallote):

i assume you know what those letters mean right?

OpenStudy (anonymous):

yes lol

OpenStudy (lgbasallote):

okay then. so what's your answer here?